OpenStudy (anonymous):

Zach borrowed $4000 from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved $5200 to pay back to his grandfather. At what interest rate did Zach borrow the money? Round to the nearest whole percent.

OpenStudy (anonymous):

First you should identify your variables and remember the formula, I=prt ( Interest, Principle, and Rate)

OpenStudy (anonymous):

The 2 years is the rate, because it shows time. The 4000 is the principle because that the amount it was borrowed. So if you're just looking for the rate, then I guess you multiply the 2 and 4000 :0 But wait!! The 2 from 2 years must be written as a decimal, so it's .02

OpenStudy (anonymous):

I think it's .02.. I can't remember exactly..

OpenStudy (anonymous):

Oh I left out a piece of information. Ignore what I just said. I'll get back to solving it

OpenStudy (anonymous):

its either 12 % 15% 18 % or 38 %

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I really don't know what to do. Sorry! I'll come back to this question later if it's still opened.. I 'm stuck on it too!

OpenStudy (anonymous):

This might help : http://www.teacherschoice.com.au/maths_library/money/simple_interest.htm

OpenStudy (anonymous):

Ok, so the way you need to do this is as how @Studybug123 started it. Your $5,200 is the Principle, the Time is 2 (I assume it's not compounding interest), and you're looking for I (the interest rate).

OpenStudy (anonymous):

Whoops - the $4,000 is the P, not the 5,200. So, can you set up that equation now?

OpenStudy (anonymous):

So I was kind of on the right track?

OpenStudy (anonymous):

Absolutely @Studybug123 - that's the right equation to use. Well done :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

ive never done these before

OpenStudy (anonymous):

Well then I guess you're gonna learn something NEW today! :)

OpenStudy (anonymous):

Ok, so the equation you need to use is I = PRT. I = the amount more than what you borrowed you pay back (hint: 5,200 - 4,000) P = original amount you borrowed R = Rate at which you borrowed (this is the one you have to solve for) T = Time (how many years was it?) So, try and set that up (don't worry we'll both help)

OpenStudy (anonymous):

wouldnt it be 5200=4000*rate*2 Solve for rate?

OpenStudy (anonymous):

That's what I kinda of thought @timo86m

OpenStudy (anonymous):

5200/(4000*2)=rate

OpenStudy (anonymous):

but there's no answer choice for that

OpenStudy (anonymous):

Hmm, so say the interest rate was 10% @timo86m wouldn't that give you this: 5200 = 4,000 x 10% x 2 So that's not correct, right?

OpenStudy (anonymous):

I believe they are asking for the interest rate as in percent, not a value of the money.. That's why I got confused

OpenStudy (anonymous):

@amberose10 - are you able to take a guess at how to set up the equation?

OpenStudy (anonymous):

@Studybug123 The I is an amount of money, the R is a percent

OpenStudy (anonymous):

Yes I know that and the answer choices ask in percent, so you have to find the Rate, right?

OpenStudy (anonymous):

the choice awsners are in percents .

OpenStudy (anonymous):

i know its 4000 x rATE X 2

OpenStudy (anonymous):

= 5200

OpenStudy (anonymous):

@amberose10 Almost - the I is the interest - so it's 5200 - 4000. The interest is the amount MORE than the principle you have to pay back, make sense?

OpenStudy (anonymous):

so 1200 ?

OpenStudy (anonymous):

its either 12 % 15% 18 % or 38 %

OpenStudy (anonymous):

4000=5200*rate*2 sfrate

OpenStudy (anonymous):

4000/(5200*2)=rate i got .38% anyone else got that?

OpenStudy (anonymous):

Oh I see. I got that too if you do it that way

OpenStudy (anonymous):

I think I got my variables confused and set up the wrong way

OpenStudy (anonymous):

thanks tim !

OpenStudy (anonymous):

@amberose10 really really learn the vocabulary the math is easy it is the vocabulary that is hard.

OpenStudy (anonymous):

thanks to everyone !

OpenStudy (anonymous):

The formula is Interest divided by principle times the time.

OpenStudy (anonymous):

got it .

OpenStudy (anonymous):

so yeah I guess just plug it in. I mixed up by variables

OpenStudy (anonymous):

Interest divided by principle times the time is to find the rate ONLY. Keep that in mind.

OpenStudy (anonymous):

http://www.teacherschoice.com.au/maths_library/money/simple_interest.htm This website gives you the 3 different formulas to find interest, rate, and time

OpenStudy (anonymous):

So, what I got was this: 1200 = 4000 x R x 2 I = PxRxT 1200 = 8000 x R 1200/8000 = R @amberose10 - Do you understand how I got that? The answer isn't 38% fyi...

OpenStudy (anonymous):

yes im following ...

OpenStudy (anonymous):

its 15 %

OpenStudy (anonymous):

@timo86m - almost, I think you just got the variables mixed up. P = 4,000 (original amount) I = 1,200 (the difference between the principal and the amount owed) - in this case 5200 - 4000 R = Rate T = Time So, in your example, it should have been 1200/8000 - 5,200 is never in the equation. Good work though!

OpenStudy (anonymous):

Yup, you got it @amberose10

OpenStudy (anonymous):

yay ! thanks !